# Policies for multiplexed quantum repeaters: theory and practical performance analysis

### v1

@@ V.2 Distillation policies
The discussion above leads us to the next question pertaining to the usefulness of distillation-based policies. We are yet to show whether there is any advantage in performing distillation. Distillation is a probabilistic process, and even when it succeeds it leads to a reduction in the number of active links. Hence, it is not manifest that a policy without distillation, such SN or FN swap-asap alone, cannot perform better than its distillation-based counterpart. Let us look at two cases: when fresh elementary links are perfect, i.e., $m_{0}=0$, and when they are imperfect, say $m_{0}=1$ for concreteness. In Fig. 10 we look at a seven-node chain with $n_{ch}=12$ channels, $p_{sw}=0.5$, and $m^{\star}=8$ ($m_{0}=1$ corresponds to roughly $90\%$ fidelity in this case). We choose FN swap-asap as our swapping policy. We show the average waiting time improvement factor for distillation-based policies, defined as the ratio of average waiting time without distillation and with distillation as a function of the elementary link success probability $p_{\ell}$. An improvement factor of greater than 1 indicates an advantage in using distillation and vice versa. We see in Fig. 10 (top) that distillation is mostly detrimental when freshly produced elementary links are perfect (100% fidelity). Both swap-distill and distill-swap increase the average waiting time, for all values of the elementary link probability. Further, the disadvantage is higher in the case of distill-swap. This can understood by the following observation. The primary role of distillation in the case when elementary links are perfect is to reduce the age of long or older links obtained as a consequence of waiting for other links and/or via entanglement swaps. The longest links tend to be the oldest ones in the network, due to the $m_{1}+m_{2}$ age-update rule for entanglement swapping. If we distill first, then initially, since most links are perfect, distillation is never invoked and after entanglement swaps, when distillation is needed, we have to wait an extra time step before distillation is attempted. This extra time step reduces the success probability of distillation of already old links, following Eq. (2). This is also the reason why swap-distill has some positive improvement in some parameter regimes. Although not shown in Fig. 10, there exists a small parameter regime of low $m^{\star},p_{\ell}$ and $n_{ch}$ in which swap-distill leads to a small positive improvement (3-5%). The choice of parameters in the figure is motivated by the intent to show the most prominent and overarching trends. On the other hand, When elementary links are imperfect, distillation is essential in increasing the scale, fidelity and throughput of the network, indicated by the significant reduction in waiting time seen in Fig. 10 (bottom). When the elementary link probability is low, distill-swap performs much better, and when it is is high, swap-distill catches up in terms of improvement. This is in agreement with the results shown in Fig. 8 and 9. Although, the trends of improvement factor variation with increasing elementary link probability change between Fig. 9 and 10. In terms of the end-to-end link fidelity swap-distill provides negligible advantage but distill-swap provides substantial advantage as shown in Fig. 10 (bottom).
V.3.1 Comparison with the SN doubling policy
Figure 11: Comparison between the SN doubling and FN and SN swap-asap policies with distillation according to distill-swap. (Top) FN swap-asap provides smaller average waiting times than SN doubling, similar to the results in Sec. III. (Bottom) The SN swap-asap policy is able to outperform the SN doubling policy when considering the average age of the youngest end-to-end link.
Let us now provide a comparison with the widely studied nested purification and doubling swapping policy. For this comparison, we fix the distillation order according to distill-swap.
In Fig. 11 (top) we first consider the average waiting time. We see that, FN swap-asap outperforms doubling, similar to the results in Sec. III in terms of average waiting time. Furthermore, when considering at average age of the youngest end-to-end link, we see in Fig. 11 (bottom) that the SN swap-asap policy is able to outperform the SN doubling policy by providing higher end-to-end fidelity. These trends are similar to those obtained in Fig. 7 in Sec. III, and further establish the superiority of swap-asap based multiplexing policies over the doubling approach in generally resource-constrained settings, and more specifically when finite coherence times for quantum memories are considered.
V.4 Main messages for distillation
@@
    Yes, when fresh elementary links are not perfect. In this case, distillation leads to a significant improvement in both the average waiting time and the average age of the end-to-end link.
    No, when fresh elementary links are perfect. In this case, distillation offers either either negligible advantage or is in fact detrimental for most parameter regimes.
    More generally, the more resource constrained the scenario, i.e., in terms of network’s hardware parameters, the more useful distillation is.
@@
    If links are generally low fidelity either due to low $p_{\ell}$ or low initial fidelity, distilling first is more advantageous, in the opposite case it is better to swap first.
@@
    More generally, in resource constrained scenarios, i.e., in terms of network’s hardware parameters, it is more useful to distill first and then swap.
VI Classical communication cost for multiplexing based swap-asap policies
In a realistic near-term entanglement distribution setting with imperfect memories, considering classical communication (CC) costs is imperative. Since entanglement distribution at the elementary link level is not deterministic, it must be heralded via classical signals. This heralding sets a natural time scale for entanglement distribution through a network. The heralding time is equal to the classical communication time between two adjacent nodes. For the purposes of our simulation, this heralding time becomes the duration of one time step. Since entanglement swapping is also often probabilistic, two distinct classes of entanglement distribution approaches can be envisaged:
  • A local approach, where all elementary link generation and entanglement swapping attempts are made agnostic of the success or failure of prior swaps. This would then mean that all elementary links (once generated successfully) are retained up to the cutoff time. When an entanglement swap has failed, resulting in both participating links to become inactive, the surrounding nodes must act as if the link is still active until it has reached its expected cutoff. All the classical communication about successful swaps is then done at the end of the protocol. Such an approach is fully local and therefore all policy decisions are taken by individual nodes, without any collaboration or exchange of information. The benefit is that no classical communication overheads exist except for heralding of elementary links, and therefore every time step has duration equal to the heralding time only. The cost is instead paid by having to wait for a link to reach its cutoff time, even though it might have become inactive long before that.
In previous works such as Refs. [42, 43], global policies were shown to optimize the average waiting time and fidelity; however, classical communication costs were not included. Furthermore, in Ref. [43], a “quasi-local” approach to entanglement distribution policies was proposed, in which there is knowledge of the network state up to some length scales only, and not globally. In this case, an advantage over fully local policies could still be obtained. We know investigate whether the advantages in waiting time and fidelity gained by global and quasi-local knowledge and collaboration can be retained when CC overheads are taken into account. The multiplexing based FN and SN swap-asap policies that we have introduced in this work are quasi-local, and they have classical communication (CC) overheads. Indeed, whenever an entanglement swapping decision is made, all the memories at the node are assumed to have knowledge of which nodes/memories they are connected to. This information is needed to decide the order of entanglement swaps based on age (SN) or length (FN) of the links. Thus, every node has some “quasi-local” knowledge of the network state but not the “global” network state. In the previous sections, the CC overheads of the FN and SN swap-asap were not taken into account when determining the waiting time. Let us now take the CC overheads into account. We limit ourselves in this section to considering the FN and SN swap-asap policies without distillation, and hence the CC costs discussed will only be the ones associated with entanglement swapping.
Below is a summary of how we have taken classical communication into account in our model.
  • In order to include the CC overheads, we approximate the CC time by adding $t_{cc}$ time steps to every network evolution time step, where $t_{cc}$ is determined by the length of the longest link involved in any entanglement swap occurring in that step. This allows enough time for the classical communication for all entanglement swaps attempted in that time step, and at the end of such a CC-accounted time step, all nodes are aware of which nodes they are connected to and the ages of those links (which can also be classically communicated). This is an upper bound to the CC cost, since in practice nodes could perform the next required action as soon as they receive the CC tagged for or relevant to them.
  • In order to benchmark our quasi-local policies fairly against fully-local policies, some modifications need to be made to the way the network evolves under the fully-local policies. These modifications are necessary, because when entanglement swaps are probabilistic, if links need to be restarted after a failed Bell state measurement, but before their actual memory cutoff, some CC will be needed. Therefore, if a policy has to be fully-local and free of any CC overheads, nodes must be agnostic to entanglement swapping failures. All nodes, once they have a heralded entangled link, must retain them until their anticipated memory cutoff age $m^{\star}$. Once this restriction is made, the only available information to nodes that can be used to decide the rank of states is the perceived or local ages of its memories. In other words, every link in the network now has two ages: one is its real age, as determined by probabilistic swaps in the evolution “in real-time”; the second is its perceived age, which is the age that the nodes that hold the link assume without any knowledge of the outcomes of entanglement swaps. When the perceived ages are used to take policy decision in the network, there is no CC overhead during the evolution. We call such a policy local age-based (ranking) multiplexed swap-asap policy. For such a policy, the CC cost is give by only one round of end-to-end communication time, in the last time step.
Figure 12: Average waiting time (top) and average age of an end-to-end link (bottom) as a function of the elementary link success probability when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy. The FN swap-asap policy provides smaller values for both figures of merit even when CC costs are included. Similar trends can also be seen for the SN swap-asap policy.
Having discussed our models for classical communication, before presenting simulation results, let us make a brief remark. In a full simulation, such as those which can be done via quantum network simulators such as NetSquid [35], QuISP [38], SeQUeNCe [36], etc., time stamps are assigned to each quantum operation and tasks are queued, such that classical communication costs are intrinsically kept track of. Our simulations focus on optimization of policies, similar in spirit to rule sets used by many of the above mentioned simulators; see also Refs. [80, 81]. The task that we undertake here is highlighting that CC costs are relevant to the performance of policies or rule sets for entanglement distribution in a realistic setting.
Let us first look at the average waiting time and average age of the end-to-end link for the FN swap-asap policy when CC costs are included, as a function of the elementary link success probability. As mentioned above, we benchmark this policy against the local age-based swap-asap policy. From Fig. 12, it is clear that substantial advantage in terms of both figures of merit is retained even when CC times are taken into consideration.
Figure 13: (Top) Average waiting time as a function of the number of nodes when classical communication times are included. We compare the quasi-local FN swap-asap policy with the fully-local age-based swap-asap policy. The FN swap-asap policy provides smaller waiting times, even when CC costs are included. More surprisingly the advantage grows with an increasing number of nodes. This provides evidence of the scalability of the advantage gained by using knowledge of the network state. (Bottom) Number distribution of lengths of entanglement swaps in the FN swap-asap policy. We observe an exponential suppression of entanglement swaps of the order of the length of the chain, which is an important reason for the retained advantage of this policy, even with classical communication included.
A more surprising observation is the fact that even as the number of nodes increases, the advantage of FN swap-asap over the local age-based ranking policy remains strong—in fact, the advantage increases with an increasing number of nodes. In Fig. 13 (top), we show the average waiting times for the local age-based ranking and the FN swap-asap policies as a function of the number of nodes. Intuitively, we might have guessed that as the number of nodes increases, so too would the CC cost, because longer and longer swaps would be required to distribute entanglement between the end points of the chain, making quasi-local policies worse off in terms of average waiting times. It is therefore surprising that not only a large improvement using the FN and SN swap-asap policies is obtained for long chains, but further that the improvement increases with the growing scale (number of nodes) of the network.
Nonetheless, we can explain this non-intuitive result as follows. First, we remind the reader that although fully-local policies have the advantage of no CC overheads, they suffer in terms of average waiting time because due to the agnostic entanglement swapping. Waiting to restart links that are already dead is detrimental to the average waiting time. Second, a quasi-local policy, as the name suggests, is not fully global, and hence CC costs in a given time-step are upper-bounded by the length of the longest entanglement swapping operation. We define the length of an entanglement swapping operation as the length (in terms of number of nodes) of the longer of two links involved in the swap. In a typical trajectory of a network, from completely disconnected to an end-to-end connected link, for the FN swap-asap policy, the number of entanglement swapping operations of length equal to the number of nodes is roughly $O(1)$, i.e., roughly constant with respect to the number of nodes. This is seen in Fig. 13 (bottom). The average number distribution of the length of entanglement swaps shows an exponentially suppressed tail (linear fall on the logarithmic scale). Therefore, time steps in a typical MDP trajectory that contribute the highest to the CC overhead are rare. This justifies and supports the use of quasi-local policies over global policies for large networks, because in the latter, the CC cost is equal for each time-step and equal to the full network size, while for quasi-local policies the CC cost can be considerably less.
VI.1 Comparison amongst quasi-local policies
Figure 14: Comparison of average waiting times for different quasi-local policies when classical communication costs are included. In the parameter regime selected for this example, reflecting a relatively resource-constrained scenario, SN swap-asap outperforms both FN swap-asap and SN doubling for the low elementary link success probability ($p_{\ell}$) regime. This could be attributed to preference given to shorter swaps by SN and hence the avoidance of large CC costs. On the other hand, when $p_{\ell}$ is larger, SN doubling performs the best, followed by SN swap-asap and then by FN swap-asap. Therefore, in this regime it is better to perform moderately longer swaps to minimize the waiting time.
We now compare the performance of different quasi-local policies when classical communication (CC) costs are included. We saw in Fig. 7 that the FN swap-asap policy provided the smallest waiting time amongst all quasi-local policies and thus also outperformed the doubling policy. Further, we observed that SN swap-asap performed worse than both SN doubling and FN swap-asap in terms of the average waiting time for end-to-end connection. With CC costs included, new interesting trends emerge in a very similar resource-constrained scenario, as shown in Fig. 14. First, in the low $p_{\ell}$ regime, we see that it is SN swap-asap that provides the smallest waiting time, followed by FN swap-asap, and then very closely followed by SN doubling. This result can be attributed to the preference given to the shortest swaps (youngest links are generally the shortest ones) by SN swap-asap compared to both the FN swap-asap and SN doubling policies. Consequently, SN swap-asap is able to avoid any large CC costs. In a scenario of low $m^{\star}$, as considered in this example, when $p_{\ell}$ is also low, most links are already very close to their cutoff time when an entanglement swap can be attempted; thus, it is essential to avoid any large CC costs, thus making SN swap-asap the policy of choice. On the other hand, when $p_{\ell}$ is larger, SN doubling performs the best, followed by SN swap-asap, and then by FN swap-asap. This result can be attributed to the fact that, in this regime, it is better to perform moderately longer swaps to minimize the waiting time. As mentioned above, we have not considered distillation-based policies here, since our model does not account for CC costs associated with distillation.
VII Proposed Experimental Implementation
@@
Figure 15: (a) Proposed implementation of multiplexing policy using a high-dimensional biphoton frequency comb. (b) Measured frequency correlation matrix of the biphoton frequency comb. (c) Polarization entanglement measurements of a biphoton frequency comb. (d) Schmidt number and Hilbert-space dimensionality scaling for different dimensions in a biphoton frequency comb.
@@
Based on the multiplexed model introduced in Fig. 1, in this section we propose a proof-of-principle experimental realization of a quantum network with multiplexing using a high-dimensional biphoton frequency comb (BFC) in Fig. 15. First, the BFC state is prepared by passing the spontaneous parametric down-conversion (SPDC) photons generated from a type-II periodically-poled KTiOPO4 (ppKTP) waveguide through a fiber Fabry–Pérot cavity (FFPC) [82, 83]. Due to the type-II phase-matching condition, the signal and idler photons can be efficiently separated by a polarization beam-splitter (PBS), enabling deterministic BFC generation without post-selection. The signal and idler photons are then temporally overlapped at a fiber beamsplitter with orthogonal polarizations, by using a tunable delay line (DL) and polarization controllers [82]. This configuration generates polarization entangled BFC state
@@
We perform the Schmidt mode decomposition of the measured joint spectral intensity [82], and summarize the extracted Schmidt mode number ($K$), and the corresponding Hilbert space dimension ($K^{2}$). When reconstructing the density matrix, we observe that the Schmidt mode number and Hilbert-space dimension scales with the dimension $d$ of the biphoton frequency comb, as shown in Fig. 15(d). The Schmidt mode number is $2.98$ with $d=3$, and $K=7.03$ when we increase the dimension to $d=9$. The deviation from ideal scaling in Fig. 15(d) mainly comes from the sinc function of the SPDC spectrum, which can be circumvented by using a flat-top broadband SPDC source. The Hilbert-space dimension can be further increased by using a SPDC source with broader spectrum bandwidth and FFPC with smaller FSR. We also note that by passing only the signal photons through a FFPC, the idler photons will still exhibit a comb-like behavior [84, 85]. Such a scheme can provide higher photon flux due to less filtering loss, and with potentially higher secure key rates for quantum key distribution [86, 87, 88]. Although we have already experimentally implemented the multiplexed source required for the proposed experimental implementation, integration of the scheme with quantum memories, implementing entanglement swapping, and distillation of links has not been done yet. At the same time, there has been great progress in each of these key steps for the construction of a long-distance quantum network.
VII.1 Calculation of elementary link probability, time step, and memory cutoff
@@
VII.2 Figures of merit and repeater chain design for existing memory platforms
@@
To answer this question, let us set the end-to-end distance to be 100 km. We also use the FN swap-asap policy for illustrative purposes, and we also include the classical communication overheads, but we leave out distillation-based policies. We also let the entanglement swapping success probability be $p_{sw}=0.5$ and the number of channels for multiplexing be $n_{ch}=9$. The other physical parameters are chosen to reflect the proposed experimental implementation above. For the quantum memories, we choose two state-of-the-art quantum memory platforms for our analysis, a rare-Earth metal based memory ($\rm{Pr}^{3+}$ ions) and a diamond vacancy based memory.
Figure 16: Average waiting time (top), including classical communication overheads, and average fidelity of the end-to-end link (bottom) for entanglement distribution along a 100 km repeater chain with rare-Earth ($\text{Pr}^{3+}$ ion) (left) and diamond vacancy (right) based memories using the FN swap-asap policy. Even though the $T_{2}$ coherence times and elementary link probabilities are relatively low, multiplexed policies provide reasonable communication rates and end-to-end link fidelities. Using five to six repeaters between the end nodes seems to be optimal for both memories in terms of waiting times.
We plot the average waiting time and fidelity in Fig. 16. We tabulate the values for the fidelity and entanglement distribution rate for the two memory platforms in the optimal setting, along with the hardware parameters, in Table 2. The optimal number of nodes in terms of the waiting time seems to be around five to seven for both memory platforms, which have $T_{2}$ times on the order of a few milliseconds. From the minimum value of the waiting time, we anticipate entanglement distribution rates ranging from a few tens of Hertz to a few hundred Hertz, with an average end-to-end link fidelity of 70-75%. At the same time, for the end-to-end fidelity, the results of the two platforms do not have the same trends. For the diamond vacancy based memory, the fidelity seems to not have its minimum at around the same point as the waiting time, indicating a direct trade-off between the two figures of merit. On the other hand, for the rare-Earth metal based memory, the fidelity decreases with an increasing number of nodes.
@@
Rare-Earth 1.2 ms 0.69 0.98 0.046 9 0 200 Hz 0.72
Diamond 13 ms 0.20 0.89 0.008 156 24 50 Hz 0.74
@@
Table 2: Hardware parameters for two quantum memory platforms (rare-earth ion and diamond vacancy), and their respective optimal entanglement distribution rates ($R$) and end-to-end fidelities ($f$), for a 100 km repeater chain, based on the data in Fig. 16. $T_{2}$ is the coherence time of the memory, $\eta$ is the absorption efficiency (Debye–Waller factor), and $f_{0}$ is the fidelity of a fresh elementary link. Also shown are the parameters $p_{\ell}$, $m^{\star}$, and $m_{0}$ of our theoretical model.
VIII Conclusions and directions for future work
In this paper, we have looked at practical multiplexing-based policies for long-distance entanglement distribution using quantum repeaters with multiple memories. We have proposed two paradigmatic policies, namely farthest neighbour (FN) swap-asap and strongest neighbour (SN) swap-asap, adapting swap-as-soon-as-possible policies for multiplexing-based linear networks. We have benchmarked these policies against a non-multiplexed, or parallel swap-asap policy, a policy that exploits multiplexing but does not use knowledge of the network state, and the widely studied doubling policy. We have shown via simulations that the FN and SN swap-asap policies can yield a considerable advantage over the alternative policies considered, both in terms of reducing the average waiting time increasing the fidelity of the end-to-end link. These advantages occur in the most relevant parameter regimes for near-term quantum networks, which correspond to the most resource-constrained settings.
We then considered policies with entanglement distillation, and we proposed a new policy that we call distill-asap. This policy combines the benefits of existing distillation policies, like the banded, greedy and pumping approaches, and thus is able to outperform the well-studied doubling policy nested with distillation. We also provide answers to two important policy questions related to distillation: When is distillation useful, and when it is useful, should we distill then swap, or the other way around? In this direction, the next question of immediate interest for future work is: “How much to distill?” Indeed, in this work, we only considered entanglement distillation policies that take $N$ links and distill them to one, but one could instead distill $N$ links to some $K$ links greater than one, and such protocols have been the study of recent work [91].

### v2

@@ V Unraveling the role of entanglement distillation policies for multiplexing
The discussion above leads us to the next question pertaining to the usefulness of distillation-based policies. We are yet to show whether there is any advantage in performing distillation. Distillation is a probabilistic process, and even when it succeeds it leads to a reduction in the number of active links. Hence, it is not obvious that a policy without distillation, such SN or FN swap-asap alone, cannot perform better than its distillation-based counterpart. Let us look at two cases: when fresh elementary links are perfect, i.e., $m_{0}=0$, and when they are imperfect, say $m_{0}=1$ for concreteness. In Fig. 10 we look at a seven-node chain with $n_{ch}=12$, $p_{sw}=0.5$, and $m^{\star}=8$ ($m_{0}=1$ corresponds to roughly $90\%$ fidelity in this case). We choose FN swap-asap as our swapping policy. We show the average waiting time improvement factor for distillation-based policies, defined as the ratio of average waiting time without distillation and with distillation as a function of the elementary link success probability $p_{\ell}$. An improvement factor of greater than 1 indicates an advantage in using distillation and vice versa. We see in Fig. 10 (top) that distillation is mostly detrimental when freshly produced elementary links are perfect (100% fidelity). Both swap-distill and distill-swap increase the average waiting time, for all values of the elementary link probability. Further, the disadvantage is higher in the case of distill-swap. This can understood by the following observation. The primary role of distillation in the case when elementary links are perfect is to reduce the age of long or older links obtained as a consequence of waiting for other links and/or via entanglement swaps. The longest links tend to be the oldest ones in the network, due to the $m_{1}+m_{2}$ age-update rule for entanglement swapping. If we distill first, then initially, since most links are perfect, distillation is never invoked and after entanglement swaps, when distillation is needed, we have to wait an extra time step before distillation is attempted. This extra time step reduces the success probability of distillation of already old links, following Eq. (2). This effect is even more magnified when CC costs are added, since swaps will now lead to even older links. This is also the reason why swap-distill has some positive improvement in some parameter regimes. Although not shown in Fig. 10, there exists a small parameter regime of low $m^{\star},p_{\ell}$ and $n_{ch}$ in which swap-distill leads to a small positive improvement (3-5%). The choice of parameters in the figure is motivated by the intent to show the most prominent and overarching trends. On the other hand, when elementary links are imperfect, distillation is essential in increasing the scale, fidelity and throughput of the network, indicated by the significant reduction in waiting time seen in Fig. 10 (bottom). When the elementary link probability is low, distill-swap performs much better, and when it is is high, swap-distill catches up in terms of improvement. This is in agreement with the results shown in Fig. 8 and 9. Although, the trends of improvement factor variation with increasing elementary link probability change between Fig. 9 and 10. In terms of the end-to-end link fidelity swap-distill provides negligible advantage but distill-swap provides substantial advantage as shown in Fig. 10 (bottom).
V.1 Main messages for distillation
@@
    Yes (generally), when fresh elementary links are not perfect. In this case, distillation leads to a significant improvement in both the average waiting time and the average age of the end-to-end link. Although this choice is not straightforward, in extremely resource constrained scenarios, such as very low $p_{\ell}$ or $p_{sw}$ it might be disadvantageous to distill (see Sec. VI.2), even when fresh links are imperfect.
    No, when fresh elementary links are perfect. In this case, distillation offers either negligible advantage or is in fact detrimental for most parameter regimes.
    More generally, the more resource constrained the scenario, i.e., in terms of network’s hardware parameters, the more useful distillation is. The observations made in this section show that detailed simulations are indispensable for analysis of the role of distillation in repeater chains.
@@
    If links are have low $p_{\ell}$ or low initial fidelity, distilling first is more advantageous, in the opposite case it is better to swap first.
@@
    More generally, in resource constrained scenarios, i.e., in terms of network’s hardware parameters, it is more useful to distill first and then swap. Further, classical communication overheads only lead to quantitative changes in these trends.
VI Experimental Implementation
@@
Figure 11: (a) Proposed implementation of a multiplexed repeater chain using a high-dimensional biphoton frequency comb (BFC). (b) Measured frequency correlation matrix of the BFC for the multiplexed elementary link that we implemented. (c) Polarization entanglement measurements of the implemented BFC. (d) Schmidt number and Hilbert-space dimensionality scaling for different dimensions in the implemented BFC.
Based on the multiplexed model introduced in Fig. 1, in this section we propose a proof-of-principle experimental realization of a quantum network with multiplexing using a high-dimensional biphoton frequency comb (BFC) in Fig. 11.
@@
As part of this proposal, we present results of an implementation a multiplexed elementary link using a high-dimensional BFC. First, the BFC state is prepared by passing the spontaneous parametric down-conversion (SPDC) photons generated from a type-II periodically-poled KTiOPO4 (ppKTP) waveguide through a fiber Fabry–Pérot cavity (FFPC) [89, 90]. Due to the type-II phase-matching condition, the signal and idler photons can be efficiently separated by a polarization beam-splitter (PBS), enabling deterministic BFC generation without post-selection. The signal and idler photons are then temporally overlapped at a fiber beamsplitter with orthogonal polarizations, by using a tunable delay line (DL) and polarization controllers [89]. This configuration generates polarization entangled BFC state
@@
We perform the Schmidt mode decomposition of the measured joint spectral intensity [89], and summarize the extracted Schmidt mode number ($K$), and the corresponding Hilbert space dimension ($K^{2}$). When reconstructing the density matrix, we observe that the Schmidt mode number and Hilbert-space dimension scales with the dimension $d$ of the biphoton frequency comb, as shown in Fig. 11(d). The Schmidt mode number is $2.98$ with $d=3$, and $K=7.03$ when we increase the dimension to $d=9$. The deviation from ideal scaling in Fig. 11(d) mainly comes from the sinc function of the SPDC spectrum, which can be circumvented by using a flat-top broadband SPDC source. The Hilbert-space dimension can be further increased by using a SPDC source with broader spectrum bandwidth and FFPC with smaller FSR. We also note that by passing only the signal photons through a FFPC, the idler photons will still exhibit a comb-like behavior [91, 92]. Such a scheme can provide higher photon flux due to less filtering loss, and with potentially higher secure key rates for quantum key distribution [93, 94, 95, 96]. Although we have already experimentally implemented the multiplexed source required for the proposed experimental implementation, integration of the scheme with quantum memories, implementing entanglement swapping, and distillation of links has not been done yet. At the same time, there has been great progress in each of these key steps for the construction of a long-distance quantum network.
VI.1 Calculation of elementary link probability, time step, and memory cutoff
@@
VI.2 Figures of merit and repeater chain design for existing memory platforms
@@
Figure 12: Average waiting time (top), including classical communication overheads, and average fidelity of the end-to-end link (bottom) for entanglement distribution along a 100 km repeater chain with nine channels with rare-Earth ($\text{Pr}^{3+}$ ion) (left) and diamond vacancy (right) based memories using the FN swap-asap policy. Even though the $T_{2}$ coherence times and elementary link probabilities are relatively low, multiplexed policies provide reasonable communication rates and end-to-end link fidelities. Using three to five repeaters between the end nodes seems to be optimal for both memories in terms of waiting times.
To answer this question, let us set the end-to-end distance to be 100 km. We also use the FN swap-asap policy for illustrative purposes, and we also include the classical communication overheads, but we leave out distillation-based policies. We consider both deterministic and non-deterministic entanglement swapping success probabilities ($p_{sw}=1.0,p_{sw}=0.5$) and set the number of channels for multiplexing to be $n_{ch}=9$. The other physical parameters are also chosen to reflect the proposed experimental implementation above. For the quantum memories, we choose two state-of-the-art quantum memory platforms for our analysis, a rare-Earth metal based memory ($\rm{Pr}^{3+}$ ions) and a diamond vacancy based memory.
We plot the average waiting time and fidelity in Fig. 12. We tabulate the values for the fidelity and entanglement distribution rate for the two memory platforms in the optimal setting, along with the hardware parameters, in Table 2. The optimal number of nodes in terms of the waiting time seems to be around five to seven for both memory platforms, which have $T_{2}$ times on the order of a few milliseconds. From the minimum value of the waiting time, we anticipate entanglement distribution rates ranging from a few tens of Hertz to a few hundred Hertz, with an average end-to-end link fidelity of 60-70%. We also see that as the swapping success probability is increased, especially close to the optimal choice of number of repeaters, there is a substantial decrease in the waiting times, although the end-to-end fidelity does not substantially improve. This is also a artifact of choosing the farthest neighbor policy which prioritizes forming long links. Also, the results of both the platforms indicate that the end-to-end link fidelity decreases with increasing number of nodes.
Finally, we look at the performance of SN doubling and distillation-based multiplexing policies for the linear repeater chain considered above. Significant differences between swap-asap and doubling policies appear for $n>5$ (four links or more). For example, for a nine-node chain using a rare-Earth metal-based memory and 50% swapping success probability, the SN doubling policy leads to around 60% higher waiting time compared to FN swap-asap. Similar improvement using the swap-asap policy can also be seen for the Diamond vacancy-based memory.
Since, the elementary links have (nearly) unit fidelity for the rare-Earth metal-based memory, distillation is detrimental as shown by the results in Fig. 10. For the diamond vacancy-based memory, we also find that for the network parameters under consideration distillation is in fact detrimental. This could be attributed to the higher CC overheads compared to the small increase in fidelity that distillation offers in this case. For example, the maximum rates available (for optimal number of repeaters) with deterministic entanglement swapping and using FN-based distill-swap policy is around 10 Hz, which is at least ten times smaller than the rates available without distillation.
@@
Rare-Earth 1.2 ms 0.69 0.98 0.046 9 0 160 Hz 0.70
Diamond 13 ms 0.20 0.89 0.004 104 24 8 Hz 0.56
@@
Table 2: Hardware parameters for two quantum memory platforms (rare-earth ion and diamond vacancy), and their respective optimal entanglement distribution rates ($R$) and end-to-end fidelities ($f$), for a 100 km repeater chain, based on the data in Fig. 12. $T_{2}$ is the coherence time of the memory, $\eta$ is the absorption efficiency (Debye–Waller factor), and $f_{0}$ is the fidelity of a fresh elementary link. Entanglement swapping is performed using linear optics hence the success probability is 50%. Also shown are the parameters $p_{\ell}$, $m^{\star}$, and $m_{0}$ of our theoretical model.
VII Conclusions and directions for future work
Near-term, resource-constrained quantum networks require the use of hardware-aware and network state-aware policies, in order to achieve high performance in terms of end-to-end waiting times and fidelities. These policies should be quasi-local, in order to reduce the impact of classical communication (CC) costs on performance. A careful assessment of the trade-off between CC costs and performance is crucial for the identification of optimal entanglement distribution policies for different hardware parameter regimes, particularly for first-generation quantum repeaters.
In this paper, we have taken steps in these directions. We have presented practical, quasi-local multiplexing-based policies for long-distance entanglement distribution using quantum repeaters with multiple memories. We call these policies farthest neighbour (FN) swap-asap and strongest neighbour (SN) swap-asap, adapting swap-as-soon-as-possible (swap-asap) policies to multiplexing-based linear networks. These policies go beyond fully local policies, such that the nodes have knowledge of states of the other nodes in the chain, but not necessarily full, global knowledge.